# Velocity effects slightly mitigating the quantumness degradation of an Unruh-DeWitt detector

### v1

@@ Abstract
In this work, we investigate the velocity effects on information degradation due to the Unruh effect in accelerated quantum systems (with finite interaction time). We consider a detector moving along a spatial trajectory within a two-dimensional plane. The quantum systems studied were: accelerated single-qubit, quantum interferometric circuit, and which-path distinguishability circuit. Thus, for non-relativistic velocity regime, we obtained analytical expressions such as transition rates, quantum coherence, visibility, distinguishability, and the complementarity relation. On the other hand, for the ultra-relativistic velocity regime, we saw that the Unruh effect is suppressed and therefore the detector does not respond in this case. Our findings revealed that velocity effects imply mitigation of information degradation, this interesting behaviors happen because of the composite effect of both velocity and acceleration. The results obtained show that the addition of the non-relativistic, transverse and constant motion of an accelerated detector can play a protective role in quantumness in systems at high accelerations, although the effects are very small.
I Introduction
@@
Hawking’s discovery hawking1975particle that black holes radiate thermally implies that they eventually evaporate. Semiclassical reasoning indicates that the final quantum state is mixed, due to the entanglement between Hawking radiation and the internal degrees of freedom that vanish upon evaporation brustein2014horizons ; brustein2015constraints ; nikolic2024semiclassical . This conclusion is difficult to avoid, yet it conflicts with the expectation that quantum theory preserves pure states. The tension between these perspectives defines the black hole information loss paradox unruh2017information ; hawking2005information .
@@
In recent years, the relativistic quantum information (RQI) has emerged as a novel research area at the intersection of general relativity and quantum information. Its primary aim is to investigate the role of relativistic effects in quantum information processing protocols martin2011relativistic ; Mann2012 . Even without considering purely quantum gravity theories, RQI covers a broad spectrum of studies with different scopes. These include the utilization of quantum probes to explore the Unruh and Hawking effects hawking1975particle ; unruh1976 ; DeWitt1980 ; wald1994quantum , as well as the formulation of quantum protocols on communication and computation Adlam2015 ; Landulfo2016 ; Martinez2015 ; Martinez2016 ; Martinez2020 ; Martinez2021 ; wu2022genuine ; Tjoa2022 ; wu2023would ; Lapponi2023 ; wu2023does ; wu2024genuinely ; wu2025can ; liu2025lorentz ; liu2025harvesting ; tang2025observational ; liu2025simulation ; li2025does ; tang2025can . Unruh’s seminal work Unruh1981Experimental demonstrated that horizons are not exclusive to gravitational systems, but they also appear in a similar way in laboratories. In particular, sonic horizons in fluids provide an experimentally accessible environment, giving rise to thermal emission similar to Hawking radiation visser1998acoustic ; gheibi2018magnetoacoustic ; tettamanti2019quantum ; dave2022hawking ; chiofalo2024dissipative .
@@
Since the last century, the interaction between quantum mechanics and general relativity has revealed conceptually interesting phenomena through the birth of quantum field theory Nambu1961Dynamical ; Dyson1949QED ; Feynman1949positrons ; Schwinger1948QED . A paradigmatic example is provided by the Fulling-Davies-Unruh effect fulling1973 ; Davies_1975 ; unruh1976 , according to which a uniformly accelerated observer in the Minkowski vacuum perceives a thermal spectrum of particles, as if immersed in a heat bath, whereas inertial observers detect no such radiation Crispino2008 . Originally formulated by Unruh in 1976 unruh1976 , this prediction highlights the fundamental role of the Rindler horizon in accelerated frames rindler1966kruskal ; rindler2012essential ; falcone2023observing , which is mathematically analogous to the event horizon of black holes birrell1984quantum ; kothawala2024noninertial . Consequently, the notion of particle number becomes observer-dependent, as distinct quantization schemes may yield inequivalent definitions of particles grove1983notes ; wald1994quantum ; lancaster2014quantum .
@@
In order to rigorously analyze this phenomenon, as well as other quantum aspects of field theory in inertial reference frames, the Unruh-DeWitt (UDW) detector was introduced unruh1976 ; DeWitt1980 . This construction constitutes a simplified yet powerful model, wherein a two-level quantum system is coupled to a quantum field. Despite its simplicity, the UDW detector faithfully captures the essential features of how distinct classes of observers interact with quantum fields. It thereby serves as a conceptual probe, enabling one to characterize and quantify the particle content perceived as a consequence of acceleration. Within this framework, one may thus operationally define particles as those excitations registered by an idealized particle detector davies1984 .
@@
A number of strategies have been put forward to probe the detection of particles associated with the Unruh effect. Among these, approaches based on quantum coherence have recently attracted considerable attention tian2012unruh ; Wang2016 ; he2018multipartite ; Nesterov2020 ; Zhang2022 ; huang2022 ; Harikrishnan2022 ; xu2023decoherence ; Pedro2024robustness ; wu2025can . In this structure, a UDW detector couples to the quantum field, absorbing quanta, thus modifying its eigenstates and leading to coherence degradation. Several recent investigations have identified conditions under which this loss of coherence is amplified. Such amplification has been reported when the vacuum is modeled as a dispersive medium barros2024dispersive , similarly when a gravitational wave background is present barros2024detecting , and when the scalar field is quadratically coupled with the detector barros2025information . In contrast, our recent results provide evidence that the mass of the scalar field can act as a protective mechanism pedro2025mitigating .
@@
Another strategy for probing this phenomenon relies on the use of quantum interferometric circuits Costa2020interferometry ; GoodingUnruh2020interferometric ; Lopes2021interference ; Pedro2024robustness ; barros2024detecting ; pedro2025mitigating . Within this framework, the particles generated due to acceleration modify the probability amplitudes that cause degradations in the visibility and interference pattern. Furthermore, interferometric configurations can be adapted to examine matter-like features through the analysis of which-path distinguishability Jaeger1995 ; Englert1996 ; miniatura2007path ; kolavr2007path ; bera2015duality . By simultaneously accounting for both wave-like and particle-like properties, one may derive the complementarity relation within this context. Such a formulation provides a natural setting for investigating how Unruh radiation influences wave-particle duality Pedro2024robustness ; pedro2025mitigating ; barros2025information .
@@
In the last century, stationary worldlines and vacuum excitation of non-inertial detectors were considered Letaw1981stationary , however, a special type of trajectory stands out where a detector moving along a spatial trajectory in a two-dimensional spatial plane has constant independent magnitudes of both the four-dimensional acceleration and a proper time derivative of the four-dimensional acceleration, so that the direction of the acceleration rotates at a constant rate around a great circle abdolrahimi2014velocity . However, it has recently been shown that superposition of accelerated linear motion and a four-velocity component of acceleration can inhibit Fisher quantum information degradation Liu2021relativistic and quantum work degradation Hao2023Quantum .
@@
Motivated by these results, this work investigates the types of effects caused by the four-velocity component in some accelerated quantum systems constructed by UDW detectors that interact for a finite time with a massless scalar field. This work is structured as follows: In Sec. II, we present the UDW theory considering velocity effects, however, we analytically show the detector response for both velocity regimes: non-relativistic and ultra-relativistic. In Sec. III, we obtained quantum coherence by considering velocity effects for an accelerated single-qubit. In Sec. IV, we show the results on wave-particle duality considering velocity effects, where we first obtain the visibility and which-path distinguishability and then calculate the complementarity relation. In Sec. V, we clarify the assumptions and limits of our model, and finally, in Sec. VI, we present our conclusions on the present work.
@@
II Unruh-DeWitt theory with velocity effects
II.1 The theoretical model
For the purposes of this study, we begin by analyzing an UDW detector unruh1976 ; DeWitt1980 , namely, a point-like system characterized by two distinct energy levels: a ground state $|g\rangle$ and an excited state $|e\rangle$. Within this framework, the detector’s proper time is denoted by $\tau$, while its trajectory is described by the worldline $x^{\mu}(\tau)=(t(\tau),\mathbf{x}(\tau))$ where $\mu$ labels the coordinates in space-time. We focus on the scenario in which the detector couples linearly to a massless scalar field $\phi[x(\tau)]$ through its monopole moment operator $\mu(\tau)$ unruh1976 ; DeWitt1980 ; birrell1984quantum ; frolov2011introduction . Accordingly, the Hamiltonian governing the linear interaction between the detector and the scalar field takes the following form:
@@ II.1 The theoretical model
with $\Omega$ being the angular transition frequency. For $\Omega>0$, the process corresponds to the absorption of a quantum, resulting in the excitation of the detector. Conversely, when $\Omega<0$, the process corresponds to the emission of a quantum, leading to the de-excitation of the detector. Moreover, $\mathcal{W}^{+}[x(\tau),x(\tau^{\prime})]$ denotes the Wightman function, defined as $\mathcal{W}^{+}[x(\tau),x(\tau^{\prime})]=\langle 0_{\mathcal{M}}|\phi[x(\tau)]\phi[x(\tau^{\prime})]|0_{\mathcal{M}}\rangle$. It is worth noting that this expression remains invariant under time translations for both inertial and uniformly accelerated trajectories. In other words, $\mathcal{W}^{+}[x(\tau),x(\tau^{\prime})]=\mathcal{W}^{+}(\tau-\tau^{\prime})$, as discussed in refs. letaw1981quantized ; padmanabhan1982general .
Recalling the property satisfied by the Wightman function [see refs. letaw1981quantized ; sriramkumar1996finite ], namely, that $\mathcal{F}^{\pm}$ are
@@
$\mathcal{R}^{\pm}(\infty)=\int_{-\infty}^{\infty}\,\mbox{\rm d}(\Delta\tau)\,\mbox{\rm e}^{\pm i\Omega\Delta\tau}\mathcal{W}^{+}(\Delta\tau).$ (8)
with $\Delta\tau=\tau-\tau^{\prime}$. It is noteworthy that the transition probability is sensitive to the derivatives of the window function. Consequently, if the detector is switched on and off abruptly, these contributions can give rise to divergent responses. To circumvent such divergences while preserving invariance under time translations for both accelerated and inertial trajectories, we employ a carefully chosen window function profile. Naturally, these considerations lead us to adopt a Gaussian window function for the UDW detector, which offers technical advantages by allowing precise control over the window width through regulation of the integrals. Accordingly, we assume the Gaussian window function $\chi(\tau)=\text{e}^{-\frac{\tau^{2}}{2T^{2}}}$ sriramkumar1996finite . By implementing this profile and restricting the interaction to a finite time interval, the resulting expression for the transition probability per unit time is
@@
II.2 Transition probability rates
Consider a UDW detector, as described in the previous section, moving in Minkowski spacetime along an unbounded spatial trajectory within a two-dimensional spatial plane. The detector follows a path characterized by a constant squared magnitude of the four-acceleration, $a^{2}=a_{\mu}a^{\mu}$, where $a^{\mu}=\mbox{\rm d}^{2}x^{\mu}/\mbox{\rm d}\tau^{2}$, as well as a constant magnitude of the timelike proper-time derivative of the four-acceleration, $(\mbox{\rm d}a_{\mu}/\mbox{\rm d}\tau)(\mbox{\rm d}a^{\mu}/\mbox{\rm d}\tau)$. Furthermore, one component of the four-velocity, $\mbox{\rm d}y/\mbox{\rm d}\tau=w$, is maintained constant. That is, the detector moves along the following worldline per unit proper time Letaw1981stationary ; abdolrahimi2014velocity :
@@ II.2 Transition probability rates
with $x^{\mu}=(t,x,y,z)$ being the Minkowski coordinates. In this path, as shown by ref. abdolrahimi2014velocity , the Wightman function corresponding to the trajectory given by Eq. (10) is given by
@@
It is worth emphasizing that we can analyze Eq. (12) for the non-relativistic velocity regime ($w\ll 1$) and for the ultra-relativistic velocity regime ($w\to\infty$). Using these regimes is essential to reveal distinct physical behaviors and allow specific analytical analyses of the detector response.
II.2.1 Non-relativistic velocity regime
Note that in the present work, the non-relativistic velocity regime ($w\ll 1$) does not refer to a classical description of motion, but rather to the relativistic expansion at low spatial velocities compared to the speed of light, so that corrections to the detector response emerge only from terms of order $w^{2}$. We now compute the transition probability rates with a finite interaction time for the non-relativistic regime. For this, the Wightman function [Eq. (12)] in leading order for $w\ll 1$ reads
$\mathcal{W}^{+}_{w}(\Delta\tau)$ $=$ $(1-2w^{2})\mathcal{W}^{+}_{0}(\Delta\tau)-\frac{a^{2}}{16\pi^{2}}\left[\sinh^{2}{\left(a\Delta\tau-2i\epsilon a\right)\left(\frac{a\Delta\tau}{4}-i\epsilon a\right)+\frac{a^{2}\Delta\tau^{2}}{4}}\right]$ (14)
$\times$ $\left[\sinh{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)}\right]^{-1}w^{2}+\mathcal{O}(w^{4}),$
and to calculate the detector transition rate, we substitute Eq. (14) into Eq. (8), calculate the contour integrals (for more information see ref. abdolrahimi2014velocity ), and defining the following dimensionless parameters $\overline{\mathcal{R}}^{\pm}=\mathcal{R}^{\pm}/\Omega$ and $\overline{a}=a/\Omega$, thus we have the excitation probability rate

### v2

# Campus Min. Petrônio Portella - Ininga, Teresina - PI, 64049-550 - Brazil

@@ Abstract
In this work, we investigate the velocity effects on information degradation due to the Unruh effect in accelerated quantum systems (with finite interaction time). We consider a detector moving along a spatial trajectory within a two-dimensional plane, this trajectory is composed of uniform acceleration along one direction, combined with a constant four-velocity component in the plane orthogonal to the acceleration. The quantum systems studied were: accelerated single-qubit, quantum interferometric circuit, and which-path distinguishability circuit. Thus, for non-relativistic velocity regime, we obtained analytical expressions such as transition rates, quantum coherence, visibility, distinguishability, and the complementarity relation. On the other hand, for the ultra-relativistic velocity regime, we saw that the Unruh effect is suppressed and therefore the detector does not respond in this case. Our findings revealed that velocity effects imply mitigation of information degradation, this interesting behaviors happen because of the composite effect of both velocity and acceleration. The results obtained show that the addition of the non-relativistic, transverse and constant motion of an accelerated detector can play a protective role in quantumness in systems at high accelerations, although the effects are very small.
arxiv: 2510.01280
1 Introduction
@@
Hawking’s discovery Hawking (1975) that black holes radiate thermally implies that they eventually evaporate. Semiclassical reasoning indicates that the final quantum state is mixed, due to the entanglement between Hawking radiation and the internal degrees of freedom that vanish upon evaporation Brustein and Medved (2014, 2015); Nikolić (2024). This conclusion is difficult to avoid, yet it conflicts with the expectation that quantum theory preserves pure states. The tension between these perspectives defines the black hole information loss paradox Unruh and Wald (2017); Hawking (2005).
@@
In recent years, the relativistic quantum information (RQI) has emerged as a novel research area at the intersection of general relativity and quantum information. Its primary aim is to investigate the role of relativistic effects in quantum information processing protocols Martín-Martínez (2011); Mann and Ralph (2012). Even without considering purely quantum gravity theories, RQI covers a broad spectrum of studies with different scopes. These include the utilization of quantum probes to explore the Unruh and Hawking effects Hawking (1975); Unruh (1976); DeWitt (1980); Wald (1994), as well as the formulation of quantum protocols on communication and computation Adlam and Kent (2015); Landulfo (2016); Pozas-Kerstjens and Martín-Martínez (2015, 2016); Simidzija et al. (2020); Tjoa and Martín-Martínez (2021); Wu and Zeng (2022); Tjoa (2022); Wu et al. (2023b); Lapponi et al. (2023); Wu et al. (2023a, 2024, 2025); Liu et al. (2025a, b); Tang et al. (2025b); Liu et al. (2025c); Li et al. (2025); Tang et al. (2025a). Unruh’s seminal work Unruh (1981) demonstrated that horizons are not exclusive to gravitational systems, but they also appear in a similar way in laboratories. In particular, sonic horizons in fluids provide an experimentally accessible environment, giving rise to thermal emission similar to Hawking radiation Visser (1998); Gheibi et al. (2018); Tettamanti et al. (2019); Dave et al. (2022); Chiofalo et al. (2024).
@@
Since the last century, the interaction between quantum mechanics and general relativity has revealed conceptually interesting phenomena through the birth of quantum field theory Nambu and Jona-Lasinio (1961); Dyson (1949); Feynman (1949); Schwinger (1948). A paradigmatic example is provided by the Fulling-Davies-Unruh effect Fulling (1973); Davies (1975); Unruh (1976), according to which a uniformly accelerated observer in the Minkowski vacuum perceives a thermal spectrum of particles, as if immersed in a heat bath, whereas inertial observers detect no such radiation Crispino et al. (2008). Originally formulated by Unruh in 1976 Unruh (1976), this prediction highlights the fundamental role of the Rindler horizon in accelerated frames Rindler (1966, 2012); Falcone and Conti (2023), which is mathematically analogous to the event horizon of black holes Birrell and Davies (1984); Kothawala (2024). Considering the principle of equivalence, similar effects have been observed in the context of the study of black holes; see some recent studies in refs. Araújo Filho et al. (2025a); Araújo Filho (2025c, a, b); Araújo Filho et al. (2025c); Heidari et al. (2025); Araújo Filho et al. (2025b). Consequently, the notion of particle number becomes observer-dependent, as distinct quantization schemes may yield inequivalent definitions of particles Grove and Ottewill (1983); Wald (1994); Lancaster and Blundell (2014).
@@
In order to rigorously analyze this phenomenon, as well as other quantum aspects of field theory in inertial reference frames, the Unruh-DeWitt (UDW) detector was introduced Unruh (1976); DeWitt (1980). This construction constitutes a simplified yet powerful model, wherein a two-level quantum system is coupled to a quantum field. Despite its simplicity, the UDW detector faithfully captures the essential features of how distinct classes of observers interact with quantum fields. It thereby serves as a conceptual probe, enabling one to characterize and quantify the particle content perceived as a consequence of acceleration. Within this framework, one may thus operationally define particles as those excitations registered by an idealized particle detector Davies (1984).
@@
A number of strategies have been put forward to probe the detection of particles associated with the Unruh effect. Among these, approaches based on quantum coherence have recently attracted considerable attention Tian and Jing (2012); Wang et al. (2016); He et al. (2018); Nesterov et al. (2020); Zhang and Liu (2022); Huang (2022); Harikrishnan et al. (2022); Xu (2023); Barros et al. (2024); Wu et al. (2025). In this structure, a UDW detector couples to the quantum field, absorbing quanta, thus modifying its eigenstates and leading to coherence degradation. Several recent investigations have identified conditions under which this loss of coherence is amplified. Such amplification has been reported when the vacuum is modeled as a dispersive medium Barros and Costa (2024b), similarly when a gravitational wave background is present Barros and Costa (2024a), and when the scalar field is quadratically coupled with the detector Barros et al. (2025a). In contrast, our recent results provide evidence that the mass of the scalar field can act as a protective mechanism Barros et al. (2025b).
@@
Another strategy for probing this phenomenon relies on the use of quantum interferometric circuits Costa et al. (2020); Gooding et al. (2020); Lopes et al. (2021); Barros et al. (2024); Barros and Costa (2024a); Barros et al. (2025b). Within this framework, the particles generated due to acceleration modify the probability amplitudes that cause degradations in the visibility and interference pattern. Furthermore, interferometric configurations can be adapted to examine matter-like features through the analysis of which-path distinguishability Jaeger et al. (1995); Englert (1996); Miniatura et al. (2007); Kolář et al. (2007); Bera et al. (2015). By simultaneously accounting for both wave-like and particle-like properties, one may derive the complementarity relation within this context. Such a formulation provides a natural setting for investigating how Unruh radiation influences wave-particle duality Barros et al. (2024, 2025b, 2025a).
@@
In the last century, stationary worldlines and vacuum excitation of non-inertial detectors were considered Letaw (1981), however, a special type of trajectory stands out where a detector moving along a spatial trajectory in a two-dimensional spatial plane has constant independent magnitudes of both the four-dimensional acceleration and a proper time derivative of the four-dimensional acceleration, so that the direction of the acceleration rotates at a constant rate around a great circle Abdolrahimi (2014). However, it has recently been shown that superposition of accelerated linear motion and a four-velocity component of acceleration can inhibit Fisher quantum information degradation Liu et al. (2021), quantum work degradation Hao et al. (2023), and the decoherence of a quantum battery quadratic environmental coupling Mukherjee et al. (2024).
@@
Motivated by these results, this work investigates the types of effects caused by the four-velocity component in some accelerated quantum systems constructed by UDW detectors that interact for a finite time with a massless scalar field. This work is structured as follows: In Sec. 2, we present the UDW theory considering velocity effects, however, we analytically show the detector response for both velocity regimes: non-relativistic and ultra-relativistic. In Sec. 3, we obtained quantum coherence by considering velocity effects for an accelerated single-qubit. In Sec. 4, we show the results on wave-particle duality considering velocity effects, where we first obtain the visibility and which-path distinguishability and then calculate the complementarity relation. In Sec. 5, we clarify the assumptions and limits of our model, and finally, in Sec. 6, we present our conclusions on the present work.
@@
2 Unruh-DeWitt theory with velocity effects
2.1 The theoretical model
For the purposes of this study, we begin by analyzing an UDW detector Unruh (1976); DeWitt (1980), namely, a point-like system characterized by two distinct energy levels: a ground state $|g\rangle$ and an excited state $|e\rangle$. Within this framework, the detector’s proper time is denoted by $\tau$, while its trajectory is described by the worldline $x^{\mu}(\tau)=(t(\tau),\mathbf{x}(\tau))$ where $\mu$ labels the coordinates in space-time. We focus on the scenario in which the detector couples linearly to a massless scalar field $\phi[x(\tau)]$ through its monopole moment operator $\mu(\tau)$ Unruh (1976); DeWitt (1980); Birrell and Davies (1984); Frolov and Zelnikov (2011). Accordingly, the Hamiltonian governing the linear interaction between the detector and the scalar field takes the following form:
@@ 2.1 The theoretical model
with $\Omega$ being the angular transition frequency. For $\Omega>0$, the process corresponds to the absorption of a quantum, resulting in the excitation of the detector. Conversely, when $\Omega<0$, the process corresponds to the emission of a quantum, leading to the de-excitation of the detector. Moreover, $\mathcal{W}^{+}[x(\tau),x(\tau^{\prime})]$ denotes the Wightman function, defined as $\mathcal{W}^{+}[x(\tau),x(\tau^{\prime})]=\langle 0_{\mathcal{M}}|\phi[x(\tau)]\phi[x(\tau^{\prime})]|0_{\mathcal{M}}\rangle$. It is worth noting that this expression remains invariant under time translations for both inertial and uniformly accelerated trajectories. In other words, $\mathcal{W}^{+}[x(\tau),x(\tau^{\prime})]=\mathcal{W}^{+}(\tau-\tau^{\prime})$, as discussed in refs. Letaw and Pfautsch (1981); Padmanabhan (1982).
Recalling the property satisfied by the Wightman function [see refs. Letaw and Pfautsch (1981); Sriramkumar and Padmanabhan (1996)], namely, that $\mathcal{F}^{\pm}$ are
@@
$\mathcal{R}^{\pm}(\infty)=\int_{-\infty}^{\infty}\,\mbox{\rm d}(\Delta\tau)\,\mbox{\rm e}^{\pm i\Omega\Delta\tau}\mathcal{W}^{+}(\Delta\tau),$ (8)
with $\Delta\tau=\tau-\tau^{\prime}$. It is noteworthy that the transition probability is sensitive to the derivatives of the window function. Consequently, if the detector is switched on and off abruptly, these contributions can give rise to divergent responses. To circumvent such divergences while preserving invariance under time translations for both accelerated and inertial trajectories, we employ a carefully chosen window function profile. Naturally, these considerations lead us to adopt a Gaussian window function for the UDW detector, which offers technical advantages by allowing precise control over the window width through regulation of the integrals. Accordingly, we assume the Gaussian window function $\chi(\tau)=\text{e}^{-\frac{\tau^{2}}{2T^{2}}}$ Sriramkumar and Padmanabhan (1996). By implementing this profile and restricting the interaction to a finite time interval, the resulting expression for the transition probability per unit time is
@@
2.2 Transition probability rates
Consider a UDW detector, as described in the previous section, moving in Minkowski spacetime along an unbounded spatial trajectory within a two-dimensional spatial plane. The detector follows a path characterized by a constant squared magnitude of the four-acceleration, $a^{2}=a_{\mu}a^{\mu}$, where $a^{\mu}=\mbox{\rm d}^{2}x^{\mu}/\mbox{\rm d}\tau^{2}$, as well as a constant magnitude of the timelike proper-time derivative of the four-acceleration, $(\mbox{\rm d}a_{\mu}/\mbox{\rm d}\tau)(\mbox{\rm d}a^{\mu}/\mbox{\rm d}\tau)$. Furthermore, one component of the four-velocity, $\mbox{\rm d}y/\mbox{\rm d}\tau=w$, is maintained constant. That is, the detector moves along the following worldline per unit proper time Letaw (1981); Abdolrahimi (2014):
@@ 2.2 Transition probability rates
with $x^{\mu}=(t,x,y,z)$ being the Minkowski coordinates. In this path, as shown by ref. Abdolrahimi (2014), the Wightman function corresponding to the trajectory given by Eq. (10) is given by
@@
It is worth emphasizing that we can analyze Eq. (12) for the non-relativistic velocity regime ($w\ll 1$) and for the ultra-relativistic velocity regime ($w\gg 1$). Using these regimes is essential to reveal distinct physical behaviors and allow specific analytical analyses of the detector response.
2.2.1 Non-relativistic velocity regime
Note that in the present work, the non-relativistic velocity regime ($w\ll 1$) does not refer to a classical description of motion, but rather to the relativistic expansion at low spatial velocities compared to the speed of light, so that corrections to the detector response emerge only from terms of order $w^{2}$. We now compute the transition probability rates with a finite interaction time for the non-relativistic regime. For this, we derive in detail in Appendix A the expansion of the Wightman function [Eq. (12)] in the principal order for $w\ll 1$, thus according to Eq. (78) we have
$\mathcal{W}_{w}^{+}(\Delta\tau)$ $\approx$ $-\frac{a^{2}}{16\pi^{2}}\Bigg\{\frac{(1-2w^{2})}{\sinh^{2}{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)}}+\Bigg[\left(\frac{\Delta\tau}{4}-i\epsilon a\right)\sinh{\left(a\Delta\tau-2i\epsilon a\right)}+$ (14)
$+$ $\frac{(a\Delta\tau)^{2}}{4}\Bigg]\frac{w^{2}}{\sinh^{4}{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)}}\Bigg\}+\mathcal{O}(w^{4}),$
and to calculate the detector transition rate, we substitute Eq. (14) into Eq. (8), calculate the contour integrals (for more information see ref. Abdolrahimi (2014)), and defining the following dimensionless parameters $\overline{\mathcal{R}}^{\pm}=\mathcal{R}^{\pm}/\Omega$ and $\overline{a}=a/\Omega$, thus we have the excitation probability rate